OpenStudy (anonymous):

A solution is made by dissolving 25.5 grams of glucose (C6H12O6) in 398 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem.

OpenStudy (jfraser):

What's the equation for solving for a DeltaT?

OpenStudy (anonymous):

Freezing Point Depression Equation: ΔT = i Kf m ΔT= freezing point depression (T_f_ (pure solvent) - T_f_ (solux)) [°C] Kf = freezing point depression constant [cryogenic constant] (depends on the properties of the solvent) [(°C*kg/mol)] m = molality (moles of solute per kilogram of solvent) [(mol/kg)] i = Van't Hoff Factor (number of ion particles per individual molecule of solute, e.g. i = 2 for NaCl, 3 for BaCl2) --- To solve: 1.) Convert grams of glucose to moles of glucose: 25.5g gluc. * (1 mol/180g gluc.) = 1.41E-1 mol gluc. Note. Molar mass gluc. = 180g/mol 2.) Convert grams of solvent to kilograms of solvent: 398 g H2O * (1kg H2O/1000g H2O) = 0.398 kg H2O 3.) Determine molality (moles solute/kilogram solvent) of solux: 1.41E-1 mol gluc./0.398 kg H2O = 3.5427E-1 mol/kg 4.) Employ Freezing Point Depression equation and determine the freezing point of the solux: ΔT = i Kf m ΔT= ? (°C) Kf = 1.86 (°C kg mol¯1) m = 3.5427E-1 (mol/kg) i = (assumed negligible) ΔΤ = (1.86)(3.5427E-1) ΔT = 6.589Ε-1°C 5.) ΔT = 6.589Ε-1°C Freezing Point Depression: http://en.wikipedia.org/wiki/Freezing-point_depression Freezing Point Depression Calculator: http://www.calctool.org/CALC/chem/substance/fp_depression Freezing Pt. Depression (2): http://www.chemteam.info/Solutions/fp-depression-problems.html

OpenStudy (anonymous):

Wouldn't 25.5g gluc. * (1 mol/180g gluc.) = equal 0.14166, not 1.41E-1 mol gluc.?

OpenStudy (jfraser):

0.141 IS equal to 1.41E-1